# Scaling Political Texts with Large Language Models: Asking a Chatbot Might Be All You Need

### v2

@@ Appendix D.2 Tweets - fine-tuned BERT classifier
Appendix D.3 Tweets - fine-tuned GloVe classifier
@@
We trained the model for up to 1,000 epochs with a $32e-$ 3 learning rate using a categorical cross entropy loss function.
Appendix D.4 Tweets - fine-tuned Naive Bayes classifier with tf-idf language representation
@@
Appendix Appendix E Additional Results - British Party Manifestos
Appendix E.1 Scatter plots of position estimates produced by LLMs and crowdworkers versus the expert survey placement
The position estimates are those used to compute the correlations reported in the body of the article (Figure 1).
Figure A2: Position of British party manifestos on the Economic policy dimension. The x-axes report the expert survey placement analyzed in Benoit et al. (2016), and the y-axes report the position estimates obtained with LLMs, fine-tuned BERT and crowd coding. The numbers next to the dots are the years of the party manifestos.
Figure A3: Position of British party manifestos on the Social policy dimension. The x-axes report the expert survey placement analyzed in Benoit et al. (2016), and the y-axes report the position estimates obtained with LLMs, fine-tuned BERT and crowd coding. The numbers next to the dots are the years of the party manifestos.
Figure A4: Position of British party manifestos on the Immigration policy dimension. The x-axes report the expert survey placement analyzed in Benoit et al. (2016), and the y-axes report the position estimates obtained with LLMs, fine-tuned BERT and crowd coding. All the party manifestos were from 2010.
Appendix E.2 Performance of position estimates obtained by including descriptions of policy dimensions in the prompts submitted to the LLMs.
Figure A5: Replication of Figure 1 except that a description of the policy dimensions and of its extremities as similar as possible to the instructions given to crowdworkers by Benoit et al. (2016) was included in the prompts submitted to the LLMs. The numbers in bold report the Pearson correlations between the benchmark (expert survey placement) and position estimates obtained with LLMs or crowdsourced position estimates reported in Benoit et al. (2016). The numbers in parentheses indicate the number of documents for which a position estimate was obtained (position estimates are missing if the LLM did not return a position score for any of the parts of manifestos submitted to it).
Appendix Appendix F Additional Results - EU legislative speeches in 10 Languages
Appendix F.1 Scatter plots that correspond to the results reported in Figure 2
Figure A6: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning in the original languages.
Figure A7: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in English.
Figure A8: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in German.
Figure A9: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Greek.
Figure A10: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Italian.
Figure A11: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Polish.
Figure A12: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. The x-axis reports the crowdsourced position estimates analyzed in Benoit et al. (2016)), and the y-axes report the position estimates obtained with LLMs. The correlations are the Pearson correlations reported in Figure 2. Positioning using the official translation in Spanish.
Appendix Appendix G Additional Results - Senators of the 117th US Congress
Appendix G.1 Correlation between the position estimates of the senators (using the baseline approach described in the body of the article) and a position estimate measure based on campaign funding.
Appendix G.2 Scaling the aggregated text of the 100 tweets in one or two prompts of up to 4,096 tokens
You will be provided with the text of a set of tweets published by a member of the US Congress. Tweets are separated by a line break followed by the string of characters <TWEET>. <TWEET> $\ll$ Text of tweet 1 $\gg$ … <TWEET> $\ll$ Text of tweet 100 $\gg$ Where does the author of these tweets stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to “NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
For most senators, the set of 100 tweets was too long to be included in a prompt of up to 4,096 tokens. Therefore, we scaled the full set of tweets in several queries to the LLM (using the same approach as for the party manifestos) and took the weighted average of the position scores (where the weights are proportional to the number of tokens) as the position estimate of a senator.
Figure A15: Replication of Figure 3 by submitting sets of tweets in prompts with up to 4,096 tokens rather than by scaling each tweet. Each dot represents a senator (‘+’: Democrats, ‘x’: Republicans, ‘o’: others). The Y-axis reports, for each senator, the position estimates of senators based on random samples of 100 of their tweets published during the Congress session. The X-axis reports the first dimension Nokken-Poole period-specific DW-NOMINATE score.
@@ Appendix G.2 Scaling the aggregated text of the 100 tweets in one or two prompts of up to 4,096 tokens
Appendix Appendix H Additional Results - Tweets published by members of the US Congress after the training cut-off date of GPT-4
@@
Table A1: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$). The last two columns report the average number of human position ratings for tweets that received a position score, and the same quantity for those for which the LLM did not return a position score. This table includes the LLMs from the body of the paper as well as a set of additional LLMs (from the previous generation). ‘Open’ LLMs can be downloaded on the user’s machine and run locally if the machine has enough video memory. ‘Local’ execution means that the model was run on our laptop after downloading the model ‘API’ execution means that the model was run for a fee on a machine managed by the provider of the model. ‘Cloud’ execution means that the model was downloaded and run via a cloud computing service (Google Colab) on a machine equipped with a TPU or a GPU. Models in .gguf format are versions of the model with compressed (‘quantized’) weights and were run locally with the llama-cpp-python package. In the names of these models,‘qX’ indicates the number of bits used to represent the model weights. We used minimal compression that would allow us to run the model on our laptop (an Apple Macbook Pro with an M1 Max processor with 64GB that can run models as large as 40GB). Meta Llama 3 8B, Llama 2 7B and Llama 2 13B were downloaded from Huggingface.co and were run locally without any compression with the mlx-lm Python package.
Appendix H.1 Distribution of position estimates produced by LLMs and other scaling methods
@@
Figure A18: Tweets published by members of the 118th US Congress: Histograms of crowdsourced position estimates for the subset (out of the total of 899 tweets) for which each scaling method produced a position estimate.
Appendix H.2 Scaling tweets on the left-right ideological spectrum by using LLMs to obtain their typicalities in the Democratic and Republican parties
@@
Appendix H.2.1 Messages submitted to LLMs.
We used the same message structure as that explained in  Appendix C except for a change to the ‘pre_text’ part of the user messages.
Appendix H.2.2 pre_text for typicality in Democratic Party.
@@
Appendix H.2.3 pre_text for typicality in Republican Party.
@@
Appendix Appendix I Predicting the position estimates given by one human respondent to the survey of tweet positions: ENO scores
@@
The ENO scores for the tweet position estimates provided by the LLMs are all higher than 1 (except for MiXtral 8X7Bq6) and go up to 5 or 6 (for GPT-4, MiXtral 8X22B and Meta Llama 3 70Bq4, over all tweets). In the latter case, this means that the position scores provided by these models are as good predictors as the average position ratings given by 5 independent human participants in our survey of Prolific participants who reside in the US.

### v3

@@ D.2 Tweets - fine-tuned BERT classifier
We used the following (standard) parameter values:
  • Batch size: 256
  • Optimizer: adam
  • Loss function: categorical_crossentropy
  • Learning rate: 2e-5
D.3 Tweets - fine-tuned GloVe classifier
@@
We trained the model for up to 1,000 epochs with a $32e-$ 3 learning rate using a categorical cross entropy loss function and a batch size of 16,384.
D.4 Tweets - fine-tuned Naive Bayes classifier with tf-idf language representation
@@
Appendix E Additional Results - Tweets published by members of the US Congress after the training cut-off date of GPT-4
@@
Table A1: Positioning tweets published by members of the US Congress after the training cut-off date of GPT-4 on the left-right ideological spectrum ($N=899$). The last two columns report the average number of human position ratings for tweets that received a position score, and the same quantity for those for which the LLM did not return a position score. This table includes the LLMs from the body of the paper as well as a set of additional LLMs (from the previous generation). ‘Open’ LLMs can be downloaded on the user’s machine and run locally if the machine has enough video memory (under more or less restrictive licenses such as Apache 2.0, the Meta Llama 3 community license agreement, or the Mistral AI Research License). ‘Local’ execution means that the model was run on our laptop after downloading the model ‘API’ execution means that the model was run for a per-token fee on a machine managed by the provider of the model. ‘Cloud’ execution means that the model was run via a cloud computing service (Google Colab or Huggingface Inference Endpoints) on a machine equipped with a TPU or GPUs (payment per hour / compute unit). Models in .gguf format are versions of the model with compressed (‘quantized’) weights and were run locally with the llama-cpp-python package. In the full names of these models,‘QX’ indicates the number of bits used to represent the model weights. We used minimal compression that would allow us to run the model on our laptop (an Apple Macbook Pro with an M1 Max processor with 64GB that can run models as large as 40GB). Llama 3.1 8B, Llama 3 8B, Llama 2 7B, Llama 2 13B, Gemma 1.1 7B, Gemma 2 9B, Aya 23 8B were downloaded from Huggingface.co and were run locally without any compression with the mlx-lm Python package. For Gemma 2 27B, we used the mlx-lm Python package to apply an 8-bit quantization convertion and to run it locally.
E.1 Distribution of position estimates produced by LLMs and other scaling methods
@@
Figure A4: Tweets published by members of the 118th US Congress: Histograms of crowdsourced position estimates for the subset (out of the total of 899 tweets) for which each scaling method produced a position estimate).
E.2 Scaling tweets on the left-right ideological spectrum by using LLMs to obtain their typicalities in the Democratic and Republican parties
@@
E.2.1 Messages submitted to LLMs.
We used the same message structure as explained in C except for a change to the ‘pre_text’ part of the user messages.
E.2.2 pre_text for typicality in Democratic Party.
@@
E.2.3 pre_text for typicality in Republican Party.
@@
E.3 Predicting the position estimates given by one human respondent to the survey of tweet positions: ENO scores
@@
The ENO scores for the tweet position estimates provided by the LLMs are all higher than 1 and go up to 5 to 7 (for GPT-4, MiXtral 8X22B and Llama 3 70Bq4, over all tweets). In the latter case, this means that the position scores provided by these models are as good predictors as the average position ratings given by 5 independent human participants in our survey of Prolific participants who reside in the US.
@@
Appendix F Additional Results - Senators of the 117th US Congress
F.1 Correlation between the position estimates of the senators (using the baseline approach described in the body of the article) and a position estimate measure based on campaign funding.
F.2 Scaling the aggregated text of the 100 tweets in one prompt
You will be provided with the text of a set of tweets published by a member of the US Congress. Tweets are separated by a line break followed by the string of characters <TWEET>. <TWEET> $\ll$ Text of tweet 1 $\gg$ … <TWEET> $\ll$ Text of tweet N $\gg$ Where does the author of these tweets stand on the ‘left’ to ‘right’ wing scale? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. If the text does not have political content, set the score to “NA”. You will only respond with a JSON object with the key Score. Do not provide explanations.
For the models we used, this resulted in a prompt that could fit the maximum context window. This means the prompt could be processed by the model in a single query
Figure A9: Replication of Figure 2 by submitting sets of tweets a single prompt. Each dot represents a senator (‘+’: Democrats, ‘x’: Republicans, ‘o’: others). The Y-axis reports, for each senator, the position estimates of senators based on random samples of 100 of their tweets published during the Congress session. The X-axis reports the first dimension Nokken-Poole period-specific DW-NOMINATE score.
Appendix G Additional Results - British Party Manifestos
G.1 Performance of position estimates obtained by including descriptions of policy dimensions in the prompts submitted to the LLMs.
Figure A10: Positioning British party manifestos on the Economic policy dimension (left to right wing scale, partial replication of figure 3). A description of the policy dimension and of its extremities as similar as possible to the instructions given to crowdworkers by Benoit et al. (2016) was included in the prompts submitted to the LLMs (see Appendix C.5).
Figure A11: Positioning British party manifestos on the Social policy dimension (liberal to conservative scale, partial replication of figure 4). A description of the policy dimension and of its extremities as similar as possible to the instructions given to crowdworkers by Benoit et al. (2016) was included in the prompts submitted to the LLMs (see Appendix C.5).
G.2 Performance of position estimates obtained by submitting the full text of manifestos in a single prompt to the LLMs.
In the body of the article, we report the results obtained by submitting each sentence of the party manifestos to the LLMs, mimicking the approach adopted by Benoit et al. (2016) with the crowds of experts and workers. This method is somewhat slow (it takes about one or two seconds to obtain a position estimate for each sentence, and sometimes longer, for large LLMs executed locally on our laptop). Therefore, we also experimented with an approach that consists in submitting the full text of the manifesto in a single prompt. Among the LLMs we have used, the only ones that have a context window long enough to accept the manifestos in one prompt are GPT-4o, GPT-4-Turbo, and MiXtral 8X22B.
We used a prompt similar to that used for obtaining position estimates of sentences, slightly adapted to the fact that we did not ask for the position of a sentence.
You will be provided with a text from a party manifesto. $\ll$ Text from the party manifesto $\gg$ Where does this text stand on the ‘left’ to ‘right’ wing scale, in terms of economic policy? Provide your response as a score between 0 and 100 where 0 means ‘Extremely left’ and 100 means ‘Extremely right’. To compute the score, start by first identifying the parts of the text that are about economic policy. Then, compute the score based on these parts of text. If the text does not refer to economic policy, return "NA". You will only respond with a JSON object with the key Score. Do not provide explanations.
The results are reported in figures A12 and A13. In terms of overall correlations, the results are similar to those obtained with the sentence-by-sentence scaling for economic policy, and not as good for social policy. The within-party correlations are not as good as with the sentence-by-sentence approach. We conjecture that this performance difference could come from the similarity between the LLM-based and human-based approaches: when both approaches rely on sentence-by-sentence coding, the results obtained by both approaches correlate more highly than if the two approaches code text of different lengths. This suggests that using LLMs in a way that mimics the procedure used with human coders might lead to a higher match with the position estimates based on human coding. Further studying this conjecture is an interesting avenue for future research.
It is also worth noting that, even though GPT-4o and MiXtral 8X22B returned position estimates for all party manifestos, GPT-4 Turbo failed to do so for two manifestos (for social policy). In ancillary analyses, we observed this failure to return position estimates for long texts in other settings and with other models as well.
Figure A12: Positioning British party manifestos on the Economic policy dimension (left to right wing scale) in a single prompt: Pearson correlations between the benchmark (expert survey placement analyzed in Benoit et al. (2016)) and position estimates obtained with LLMs or crowdsourced position estimates reported in Benoit et al. (2016).
Figure A13: Positioning British party manifestos on the Social policy dimension (liberal to conservative scale) in a single prompt: Pearson correlations between the benchmark (expert survey placement analyzed in Benoit et al. (2016)) and position estimates obtained with LLMs or crowdsourced position estimates reported in Benoit et al. (2016).
Appendix H Additional Results - EU legislative speeches in 10 Languages
H.1 Result summary
crowd English German Greek Italian Polish Spanish crowd English 1.00 0.93 0.94 0.95 0.96 0.94 German 1.00 0.94 0.92 0.92 0.92 Greek 1.00 0.92 0.94 0.95 Italian 1.00 0.96 0.94 Polish 1.00 0.94 Spanish 1.00                   Refer to caption Refer to caption
Figure A14: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension: Pearson correlations between the benchmark (crowdsourced position estimates reported in Benoit et al. (2016)). The numbers in parentheses indicate the number of documents for which a position was obtained. As a performance benchmark, we report the pairwise correlations between crowdsourced estimates obtained from the translations of speeches in the different languages (upper left panel). These correlations are of similar magnitude.
H.2 Scatter plots for scaling of translations of speeches by LLMs and crowd workers
Figure A15: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. Positioning using the official translation in English.
Figure A16: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. Positioning using the official translation in German.
Figure A17: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. Positioning using the official translation in Greek.
Figure A18: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. Positioning using the official translation in Italian.
Figure A19: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. Positioning using the official translation in Polish.
Figure A20: Positioning EU legislative speeches in 10 languages on the ‘anti-subsidy’ to ‘pro-subsidy’ dimension. Positioning using the official translation in Spanish.